OpenStudy (anonymous):

Jack and Jazz weighs in the ratio of 5:6. After 10 years their weight will be 6:7. What are their Present ages?

OpenStudy (mathstudent55):

Not enough info. We know the ratio of their weights and the future ratio of their weights. We know nothing about their ages.

OpenStudy (anonymous):

it is very easy

OpenStudy (anonymous):

LilySwan your answer is wrong

OpenStudy (mathstudent55):

Jack weighed 100 lb. He went on a diet and lost 10 lb. How old is Jack now?

OpenStudy (anonymous):

first tell my answer .

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it has a proper data

OpenStudy (mathstudent55):

@Sachit Are the ratios age ratios, not weight ratios?

OpenStudy (anonymous):

sorry it is weight ratio

OpenStudy (mathstudent55):

And the question is still how old they are? And the given question pertains to 10 years from now?

OpenStudy (anonymous):

i am writting the question again

OpenStudy (mathstudent55):

ok

OpenStudy (anonymous):

Jack and James ages are in the ratio of 5:6. After 10 years their age will be in the ratio of 6:7. What are their Present ages?

OpenStudy (mathstudent55):

Ok. Now everything relates to ages, so we can solve it.

OpenStudy (anonymous):

now it is a right question

OpenStudy (mathstudent55):

Let's say Jack is x years old now. Also, we'll say James is y years old now. Since the ratio of their ages is 5:6, we can write an equation: x/y = 5/6 Ok so far?

OpenStudy (mathstudent55):

In 10 years, each one is 10 years older. In 10 years, Jack will be x + 10 years old. In 10 years, James will be y + 10 years old. In 10 years, the ratio of their ages is: (x + 10)/(y + 10) = 6/7

OpenStudy (mathstudent55):

Now we have a system of equations. We want to solve for x and y, the present ages.

OpenStudy (mathstudent55):